OpenStudy (anonymous):

An electron travels 1.64 m in 4.32 × 10^−8 s. How fast is it traveling? Answer in units of m/s

OpenStudy (theeric):

Hi! You'll want to remember the velocity formula, which is displacement per time. It's sort of like speed: 50 miles per hour or 70 kilometers per hour. Speed is the distance you cover in an interval of time. In the math, velocity is \[v=\frac{d}{t}\]Where \(v\) is your velocity, \(d\) is the displacement made, and \(t\) is the amount of time it too to do that. For you, your displacement is \(1.64\ [m]\) and your time interval is \(4.32\times 10^{-8}\ [s]\). So your velocity is \(1.64\ [m]\) per \(4.32\times 10^{-8}\ [s]\), or \(\Large\frac{\(1.64\ [m]}{4.32\times 10^{-8}\ [s]}\). You plug it into a calculator, and you have your answer in \([m/s]\) like the question asks for! Any questions?

OpenStudy (anonymous):

Nope got it! Thanks so much!

OpenStudy (theeric):

Cool! You're very welcome :)